OpenStudy (anonymous):

How do I graph 2x – y = 2

OpenStudy (mimi_x3):

\[2x - y =2\] y-int: x= 0 x-int: y = 0

OpenStudy (mimi_x3):

are you able to do it now?

OpenStudy (mimi_x3):

y-int: x=0 => 2(0) - y = 2 => 0 -y = 2 => -y = 2 => y = -2

OpenStudy (mimi_x3):

you try the next one :)

OpenStudy (anonymous):

yea let me try one sec

OpenStudy (anonymous):

attached

OpenStudy (mimi_x3):

so you sketched it? doesnt look right

OpenStudy (anonymous):

2x-y =2 ...add y to each side of the equation...2x= 2+y........lift 2 from each side...y=2x-2...so when x=2 then y = 2 ..and when x=3 y is 4...plot thes points on a grid and join them up...good luck